# A Survey of Robotic Language Grounding: Tradeoffs Between Symbols and Embeddings

### v1

@@ 4.4 Generalization and Bias
Many existing works that use foundation models also leverage symbols but do not speak to the symbols’ role in building practical systems. We observe that in many recent RoboNLP papers including  Ichter et al., 2022; Jiang et al., 2023; Brohan et al., 2023; Zitkovich et al., 2023; Padalkar et al., 2023 all leverage symbolic representations, either in the form of discretized action spaces or predefined skills. We would like to see more authors acknowledge that their work is made better because of the symbolic representation their model uses or propose how other symbolic representations will make the model better.
In the future, we expect models to better represent the world through these finer-grained symbolic representations. After a certain point, combining text, audio, color images, depth images, point clouds, etc., becomes functionally continuous. While the POMDP and PDDL methods are theoretically sound, in practice, they lack representational capacity outside of their domain due to the brittleness of out-of-domain representation. We believe that future successful methods will leverage deep learning methods with smarter representations of the input and output data via transformations of the intermediate representations. This includes representing color images as point clouds, maps, or meshes as derivative intermediates. We can then augment these with state representations of the environment as text or other labels. There is an eventuality where the bitter lesson (Sutton, 2019) takes over, and throwing even more parameters and data at the problem will win out, but that is predicated on an evergrowing data supply. It remains to be seen where additional data will come from and if methods that leverage only partial observations from one modality can synergize multiple modalities simultaneously.
@@ 4.5 Limitations of Natural Langauge
As robot operators, we want to be able to express the state of our robotic system in a way that is auditable by us and useful to the robot in accomplishing tasks. Many of the methods we discussed also focus on English as the representation of natural language encoded in LLMs and multimodal models. While English works well for much of the world, it does not precisely describe objects’ physical locality. Saying that an apple is to the left of an orange can put that apple in various positions with just a language description alone. While adding geometric priors could enhance the specificity of the language, we should, as a field, look to building benchmarks to evaluate the physical plausibility of these language models as they operate in different languages (Whorf, 1956; Lucy, 1992). Robot morphology and human morphology are very different. It is likely that expressing tasks as being completed by one’s arm, such as “clean the dishes with your left hand” may be completely moot to a robot that uses soft-body mechanics to interact with the world. Human language evolved in a biological domain that necessitates a more survival vocabulary. These actions are not necessarily helpful for describing robot actions and could contribute to skewing the success rates of robots in accomplishing tasks.
@@ 4.6 Safety and Interpretability
Interpretable and explainable robotic systems provide transparency in decision-making and gain trust in human-robot interaction (Gunning and Aha, 2019; Anjomshoae et al., 2019; Silva et al., 2023). Moreover, verifiable safe operation is essential for deployments that satisfy worldwide standards such as ISO 61508 (International Organization for Standardization/International Electrotechnical Commission, 2010), which defines standards for safely deploying robots in industrial factory environments worldwide. These standards require that robotic systems be mathematically proven to have a failure rate lower than $10^{-5}$ dangerous failures per hour. Robots should also provide feedback to unsatisfiable task specifications (Raman et al., 2013) or explain their actions when execution fails (Das et al., 2021). Safety can be broken down into semantic and kinematic/physical safety (Varley et al., 2024). Examples of semantic safety are never entering the nursery and always transporting a cup full of coffee in an upright orientation. Examples of kinematic safety are avoiding collisions with humans and objects and avoiding reaching joint/velocity/torque limits. Formal approaches using LLMs mostly focus on enforcing semantic safety. Many works on formal methods and safe control do not use LLMs and encode kinematic safety in trajectory optimization (Dawson et al., 2023). For example, linear temporal logic (LTL) has been used extensively to develop provably safe controllers (Lignos et al., 2015; Chinchali et al., 2012). Given a logical specification and the system model, formal methods can synthesize correct-by-construction robot controllers or provide counterexamples to explain when the task is unsatisfiable with respect to how well the model captures reality (Kress-Gazit et al., 2018). A key challenge for any safety framework of this kind is grounding predicates in real-world noisy and partial perceptual data.
Existing deep learning approaches do not have a clear pathway for achieving this level of interpretability and safety. If we look at human behavior, defining a set of safety guidelines to assess adherence has been challenging. While we can continue to apply formal representations to these models to varying levels of success, there remains a need for a system that, at a meta-level, operates on symbols but whose abilities are as general and flexible as humanity’s. For example, one of our recent papers uses a combined approach with LLMs and a formal representation, showing that we can exploit the generalization ability of LLMs with the formal safety guarantees provided by LTL to create a safe yet robust and flexible system for following commands (Yang et al., 2024). Yet this approach only scratches the surface. Much more needs to be done to integrate these systems, especially to study the interpretability and safety guarantees inherent in perceptual systems and provide formal guarantees and bounds about the behavior of deep networks.
@@ 5 Conclusion
Our review characterizes the literature in robotic language grounding along a spectrum from using more formal, biased, discrete representations to less formal, less biased, higher-dimensional continuous representations. There are benefits and tradeoffs to each approach. More formal methods induce structure that can limit the size of the learning problem and provide interpretability and formal safety guarantees. However, they also constrain the output space, limiting the flexibility and expressive power of what can be learned. Less formal methods impose fewer constraints but require more data and possibly more structured neural networks to be learned. Methods such as SayCan (Ichter et al., 2022), traditionally considered less structured, use a formal representation of the robot’s skills. Seen in this context, it is clear that a limitation for all methods is the lack of physical capability of existing robots: a key area of future work is to enable robots to perform a larger variety of tasks in a larger variety of environments. Hence, they can physically perform the tasks people ask them to do.
6 Acknowledgments
The work done by Vanya Cohen and Raymond Mooney is supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0007. The work done by Jason Xinyu Liu and Stefanie Tellex is supported by the Office of Naval Research (ONR) under grant number N00014-22-1-2592 and with support from Amazon Robotics. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies. The authors would like to thank Ankit Shah and Jessica Hodgins for their insightful feedback.

### v2

@@ 4.4 Generalization and Bias
Many existing works that use foundation models also leverage symbols but do not speak to the symbols’ role in building practical systems. We observe that in many recent RoboNLP papers including  Ahn et al. (2022); Jiang et al. (2023); Brohan et al. (2023a, b); O’Neill et al. (2024) all leverage symbolic representations, either in the form of discretized action spaces or predefined skills. We would like to see more authors acknowledge that their work is made better because of the symbolic representation their model uses or propose how other symbolic representations can improve the model.
In the future, we expect models to better represent the world through these finer-grained symbolic representations. After a certain point, combining text, audio, color images, depth images, point clouds, etc., becomes functionally continuous. While the POMDP and PDDL methods are theoretically sound, in practice, they lack representational capacity outside of their domain due to the brittleness of out-of-domain representation. We believe that future successful methods will leverage deep learning methods with smarter representations of the input and output data via transformations of the intermediate representations. This includes representing color images as point clouds, maps, or meshes as derivative intermediates. We can then augment these with state representations of the environment as text or other labels. There is an eventuality where the bitter lesson Sutton (2019) takes over, and throwing even more parameters and data at the problem will win out, but that is predicated on an evergrowing data supply. It remains to be seen where additional data will come from and if methods that leverage only partial observations from one modality can synergize multiple modalities simultaneously.
@@ 4.5 Limitations of Natural Langauge
As robot operators, we want to be able to express the state of our robotic system in a way that is auditable by us and useful to the robot in accomplishing tasks. Many of the methods we discussed also focus on English as the representation of natural language encoded in LLMs and multimodal models. While English works well for much of the world, it does not precisely describe objects’ physical locality. Saying that an apple is to the left of an orange can put that apple in various positions with just a language description alone. While adding geometric priors could enhance the specificity of the language, we should, as a field, look to building benchmarks to evaluate the physical plausibility of these language models as they operate in different languages Whorf (1956); Lucy (1992). Robot morphology and human morphology are very different. It is likely that expressing tasks as being completed by one’s arm, such as “clean the dishes with your left hand” may be completely moot to a robot that uses soft-body mechanics to interact with the world. Human language evolved in a biological domain that necessitates a more survival vocabulary. These actions are not necessarily helpful for describing robot actions and could contribute to skewing the success rates of robots in accomplishing tasks.
@@ 4.6 Safety and Interpretability
Interpretable and explainable robotic systems provide transparency in decision-making and gain trust in human-robot interaction Gunning and Aha (2019); Anjomshoae et al. (2019); Silva et al. (2023). Moreover, verifiable safe operation is essential for deployments that satisfy worldwide standards such as ISO 61508 International Organization for Standardization (2010), which defines standards for safely deploying robots in industrial factory environments worldwide. These standards require that robotic systems be mathematically proven to have a failure rate lower than $10^{-5}$ dangerous failures per hour. Robots should also provide feedback to unsatisfiable task specifications Raman et al. (2013) or explain their actions when execution fails Das et al. (2021). Safety can be classified into semantic and kinematic/physical safety Varley et al. (2024). Examples of semantic safety are never entering the nursery and always transporting a cup full of coffee in an upright orientation. Examples of kinematic safety are avoiding collisions with humans and objects and avoiding reaching joint/velocity/torque limits. Formal approaches using LLMs mostly focus on enforcing semantic safety. Many works on formal methods and safe control do not use LLMs and encode kinematic safety in trajectory optimization Dawson et al. (2023). For example, linear temporal logic (LTL) has been used extensively to develop provably safe controllers Lignos et al. (2015); Chinchali et al. (2012). Given a logical specification and the system model, formal methods can synthesize correct-by-construction robot controllers or provide counterexamples to explain when the task is unsatisfiable with respect to how well the model captures reality Kress-Gazit et al. (2018). A key challenge for any safety framework of this kind is grounding predicates in real-world noisy and partial perceptual data.
Existing deep learning approaches do not have a clear pathway for achieving this level of interpretability and safety. If we look at human behavior, defining a set of safety guidelines to assess adherence has been challenging. While we can continue to apply formal representations to these models to varying levels of success, there remains a need for a system that, at a meta-level, operates on symbols but whose abilities are as general and flexible as humanity’s. For example, one of our recent papers uses a combined approach with LLMs and a formal representation, showing that we can exploit the generalization ability of LLMs with the formal safety guarantees provided by LTL to create a safe yet robust and flexible system for following commands Yang et al. (2024). Yet this approach only scratches the surface. Much more needs to be done to integrate these systems, especially to study the interpretability and safety guarantees inherent in perceptual systems and provide formal guarantees and bounds about the behavior of deep networks.
@@ 5 Conclusion
Our review characterizes the literature in robotic language grounding along a spectrum from using more formal, biased, discrete representations to less formal, less biased, higher-dimensional continuous representations. There are benefits and tradeoffs to each approach. More formal methods induce structure that can limit the size of the learning problem and provide interpretability and formal safety guarantees. However, they also constrain the output space, limiting the flexibility and expressive power of what can be learned. Less formal methods impose fewer constraints but require more data and possibly more structured neural networks to be learned. Methods such as SayCan Ahn et al. (2022), traditionally considered less structured, use a formal representation of the robot’s skills. Seen in this context, it is clear that a limitation for all methods is the lack of physical capability of existing robots: a key area of future work is to enable robots to perform a larger variety of tasks in a larger variety of environments. Hence, they can physically perform the tasks people ask them to do.
Acknowledgments
The work done by Vanya Cohen and Raymond Mooney is supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0007. The work done by Jason Xinyu Liu and Stefanie Tellex is supported by the Office of Naval Research (ONR) under grant number N00014-22-1-2592 and with support from Amazon Robotics. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies. The authors thank Ankit Shah and Jessica Hodgins for their insightful feedback.
Contribution Statement
All authors contributed equally. Listed in alphabetical order.